• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mindy

    Log entry time 16:24:54 on June12,2010

    Entry number 326068

    keyword=06/12/10 Day Shift Summary

    8:00 AM doing dipole scan test
    10:00 AM run plan meeting
    11:30 AM end the dipole scan test
    12:00 PM start the beam recovery
    13:30 PM start C12 transverse polarization measurement
    14:00 PM MCC called: target vacuum becomes unstable
    14:30 PM from 14:00 PM, the beam ramp down to zero, and then we ask MCC to ramp up the beam 10uA per couple of minutes
    Megan called MCC shift screws to tune the Compton beam and the laser phase. In addition, Abdurahim helped check the Compton cavity power at home
    14:40 PM Compton run starts
    15:04 PM C12 production run 4476 (offline-analysis done)
    15:33 PM C12 production run 4477 (offline-analysis done)